# Sound Event Detection and Localization with Distance Estimation

### v1

@@ I Introduction
Computational Auditory Scene Analysis (CASA) has emerged as a prominent area of study in recent years [1]. The automated examination of audio content holds substantial potential for diverse practical applications, including speech recognition [2], autonomous robots [3], surveillance systems [4], and support systems for the hearing-impaired [5, 6]. CASA encompasses a spectrum of audio-related tasks, ranging from acoustic scene classification [7] and audio tagging [8] to sound source localization [9] and sound event detection [10]. Most tasks are currently approached by modelling Deep Neural Networks (DNNs).
@@ II-A Features
For Ambisonics, we accumulate the magnitude spectrograms from 4 channels into $F=64$ mel energies. In order to explore spatial cues, we extract 3 intensity vector matrices, accumulated along the mel frequencies to fit the required number of features. The overall input matrix sums up to $CH=7$ feature channels.
@@ II-B Model
Figure 1: Architecture of the deep model.
@@
The output of the model consists of $Q$ branches. Each branch is composed of two fully connected (FC) layers, the former containing 128 neurons, and the latter outputting $C_{q}$ values. The parameters of the last layers are determined by the utilized method of including SDE within the 3D SELD framework. Here, we propose two techniques:
@@
where $\mathcal{L}_{1}(\cdot)$ and $\mathcal{L}_{2}(\cdot)$ are the loss functions of the first and second branch.
We note that the classwise output of the MT approach does not allow for differentiating between overlapping sound sources of the same class, whereas the multi-ACCDDOA approach overcomes this problem. The output parameters are summarized in Table II. Models are implemented in PyTorch [23] and trained using the Adam optimizer for 250 epochs with 75 epochs of patience.
@@ III-C Results
As can be seen from most metrics, the performance of the models are significantly worse for the binaural dataset than for the Ambisonic dataset. For the multi-task approach, the error rate went up from the range of [0.63, 0.65] to [0.82, 0.87], whereas the $F_{1}$ score decrased from over 40% to the range of [16.5, 20.0]%. A similarly large effect can be seen for the localization metrics, where the DOA error ranges from 38.1° to 41.1° and the recall from 36.7% to 45.6%. This is a partly expected effect. Binaural recordings consist of only two channels as compared with four channels for FOA. Moreover, the cone of confusion effect and high directivity of the ears might impact the performance significantly. These problems can be largely overcome by utilizing a moving receiver as shown in [21, 22]. Interestingly, distance estimation has been affected to a much lesser extent than the other tasks. For the multi-task approach, the distance error stays in the range between 0.98m to 1.04m, which is roughly a 10% increase from the Ambisonic results. These results indicate that efficient distance estimation can be achieved even with binaural recordings, which shows a potential for future research.
As can be seen for both audio formats, for the multi-task approach most differences between loss functions appear for the distance error. This is an expected effect given the ACCDOA branch uses the same loss for all scenarios. For MSE and MAE, the Ambisonic model achieves a similar distance error of 0.95 m and 0.93 m, respectively. This is a reasonable performance compared with other metrics. Introducing the distance scaling brings the error further down, to 0.89m for MSPE and 0.88m for MAPE. Similarly, the binaural model achieves the lowest distance error of 0.98 m for MAPE. The SELD metrics show very limited influence of the distance loss on the training of the rest of the multi-task model. For FOA data, all scenarios achieve a fairly similar error rate between 0.63 and 0.65 and an $F_{1}$ score between 41.4% for MSE and 44.1% for MSPE. Interestingly, a higher impact can be seen for the localization metrics. Despite the DOA error staying between 21.6° and 23.2°, the recall varies between different scenarios. Compared with 61.0% for MSE, the MAE loss results in a recall of 41.1%, showing a decreased efficiency in estimated the correct number of sources. Also for binaural data, the MAE loss decreases the recall from 45.6% to 36.7%. However, for FOA the distance-scaled losses visibly improve the localization recall to 64.7% and 64.5% for MSPE and MAPE, respectively.

### v2

@@ I Introduction
Computational Auditory Scene Analysis (CASA) has emerged as a prominent area of study in recent years [1]. The automated examination of audio content holds substantial potential for diverse practical applications, including speech recognition [2], autonomous robots [3], surveillance systems [4], and support systems for the hearing-impaired [5]. CASA encompasses a spectrum of audio-related tasks, ranging from acoustic scene classification [6] and audio tagging [7] to sound source localization [8] and sound event detection [9]. Most tasks are currently approached by Deep Neural Network (DNN) models.
@@ II-A Features
For Ambisonics, we accumulate the magnitude spectrograms from 4 channels into $F=64$ mel energies. In order to explore spatial cues, we extract 3 intensity vector matrices as in [21], accumulated along the mel frequencies to fit the required number of features. The overall input matrix sums up to $CH=7$ feature channels.
@@ II-B Model
Figure 1: Architecture of the deep neural network.
@@
The output of the model consists of $Q$ branches, where $Q$ depends on the number of performed tasks. Each branch is composed of two fully connected (FC) layers, the former containing 128 neurons, and the latter outputting $C_{q}$ values. The parameters of the last layers are determined by the utilized method of including SDE within the 3D SELD framework. Here, we propose two techniques:
@@
where $\mathcal{L}_{1}(\cdot)$ and $\mathcal{L}_{2}(\cdot)$ are the losses of the both branches.
We note that the output of the MT approach does not allow for differentiating between overlapping sources of the same class, whereas the multi-ACCDDOA approach overcomes this problem. The output parameters are summarized in Table II. Models are implemented in PyTorch [23] and trained using the Adam optimizer for 250 epochs with 75 epochs of patience.
@@ III-C Results
As can be seen from most metrics, the performance of the models are significantly worse for the binaural dataset than for the Ambisonic dataset. For the multi-task approach, the error rate went up from the range of [0.63, 0.65] to [0.82, 0.87], whereas the $F_{1}$ score decrased from over 40% to the range of [16.5, 20.0]%. A similarly large effect can be seen for the localization metrics, where the DOA error ranges from 38.1° to 41.1° and the recall from 36.7% to 45.6%. This is a partly expected effect. Binaural recordings consist of only two channels as compared with four channels for FOA. Moreover, the cone of confusion effect and high directivity of the ears might impact the performance significantly. These problems can be largely overcome by utilizing a moving receiver as shown in [20, 22]. Interestingly, distance estimation has been affected to a much lesser extent than other tasks. For the MT approach, the distance error stays in the range between 0.98m to 1.04m, which is roughly a 10% increase from the Ambisonic results. These results indicate that efficient distance estimation can be achieved even with binaural recordings, which shows a potential for future research.
As can be seen for both audio formats, for the multi-task approach most differences between loss functions appear for the distance error. This is expected given the ACCDOA branch uses the same loss for all scenarios. For MSE and MAE, the Ambisonic model achieves a similar distance error of 0.95 m and 0.93 m, respectively. This is a reasonable performance compared with other metrics. Introducing the distance scaling brings the error further down, to 0.89m for MSPE and 0.88m for MAPE. Similarly, the binaural model achieves the lowest distance error of 0.98 m for MAPE. The SELD metrics show very limited influence of the distance loss on the training of the rest of the multi-task model. For FOA data, all scenarios achieve a fairly similar error rate between 0.63 and 0.65 and an $F_{1}$ score between 41.4% for MSE and 44.1% for MSPE. Interestingly, a higher impact can be seen for the localization metrics. Despite the DOA error staying between 21.6° and 23.2°, the recall varies between different scenarios. Compared with 61.0% for MSE, the MAE loss results in a recall of 41.1%, showing a decreased efficiency in estimated the correct number of sources. Also for binaural data, the MAE loss decreases the recall from 45.6% to 36.7%. However, for FOA the distance-scaled losses visibly improve the localization recall to 64.7% and 64.5% for MSPE and MAPE, respectively.